# Sharp Self-Normalized Concentration Inequalities of Marginal Mean with Sample Variance Only

### v2

@@ A.2 Proof of Results in Section 2
Proof of Theorem 1 We only focus on proving right tail of (8).
Step 1(Construction of super-martingale) Some simple algebra shows that, when $\lambda\in[0,\frac{1}{b}]$,
$e^{\lambda Z_{i}-\frac{\lambda^{2}}{2}Z_{i}^{2}}\leq 1+\lambda Z_{i}$ (A.9)
holds for any given $i$ and every $\omega\in\Omega$. Since $(Z_{i},\mathcal{A}_{i})_{i\geq 1}$ is a MDS, (A.9) implies
$\mathbb{E}\Big[\exp\Big(\lambda Z_{i}-\frac{\lambda^{2}Z_{i}^{2}}{2}\Big)|\mathcal{A}_{i-1}\Big]\leq 1+\lambda\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=1\ \$(A.10)
holds almost surely for every $\lambda\in[0,1/b]$. For this arbitrary given $\lambda\in[0,1/b]$, we also define
$L_{m,\lambda}:=\exp\Big(\lambda\mathbf{M}_{m}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}\Big)$ (A.11)
Then, using (A.10) and basic properties of MDS yields that, for any given $m\geq 1$,
$\mathbb{E}[L_{m,\lambda}|\mathcal{A}_{m-1}]=\mathbb{E}\Big[\exp\Big(\lambda\mathbf{M}_{m}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}\Big)\Big|\mathcal{A}_{m-1}\Big]$
$=$ $\exp\Big(\lambda\mathbf{M}_{m-1}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m-1}\Big)\mathbb{E}\Big[\exp\Big(\lambda W_{m}-\frac{\lambda^{2}}{2}W_{m}^{2}\Big)\Big|\mathcal{A}_{m-1}\Big]$
$\leq$ $\exp\Big(\lambda\mathbf{M}_{m-1}-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m-1}\Big)=L_{m-1,\lambda}.$
Thus, we manage to prove that, for any given $\lambda\in[0,1/b]$, $\{L_{m,\lambda},\mathcal{A}_{m}\}_{m\geq 1}$ is a super martingale Step 2 (Extension of filter and randomization of $\lambda$) Define
$L_{m}(\lambda,\cdot):=\exp\Big(\lambda\mathbf{M}_{m}(\cdot)-\frac{\lambda^{2}}{2}[\mathbf{M}]_{m}(\cdot)\Big)\ \ \text{and}\ \ g_{m}(\lambda,\omega):=\mathbb{E}[L_{m}(\lambda,\cdot)|\mathcal{A}_{m-1}](\omega),$
we have
  • (i)
    For each $m\geq 1$, both $g_{m}(\lambda,\omega)$ and $L_{m}(\lambda,\cdot)$ are real-valued and non-negative measurable mappings defined on $0,1/b]\times\Omega$;
  • (ii)
    For each given $\lambda\in(0,1/b]$, $\mathbb{P}(\omega\in\{g_{m}(\lambda,\omega)\leq L_{m-1}(\lambda,\omega)\})=1$.
Then, for any random variable $\Lambda$ taking value in $[0,1/b]$, by denoting $\mathcal{G}_{m}:=\sigma(\Lambda)\lor\mathcal{A}_{m}$ and $L_{m,\Lambda}(\cdot):=L_{m}(\Lambda(\cdot),\cdot)$, $\{L_{m,\Lambda}\}_{m\geq 1}$ is adapted to filter $\{\mathcal{G}_{m}\}_{m\geq 0}$. $\mathcal{G}_{0}$ is any sigma algebra contained by the intersection of $\bigcap_{i\geq 1}\mathcal{G}_{t}$ and independent of $\sigma(\bigcup_{i\geq 1}\mathcal{G}_{i})$. This can always be done by properly enlarging the original probability space, like introducing product space. Meanwhile, Lemma A.4 implies that
$\mathbb{E}[L_{m,\Lambda}|\mathcal{G}_{m-1}](\omega)=\mathbb{E}[L_{m}(\Lambda(\cdot),\cdot)|\mathcal{G}_{m-1}](\omega)=g_{m}(\Lambda(\omega),\omega)\ \ \text{a.s.}$ (A.12)
holds for any given $m\geq 1$. Together with Point (ii) above, we thus obtain
$\mathbb{E}[L_{m,\Lambda}|\mathcal{G}_{m-1}](\omega)=g_{m}(\Lambda(\omega),\omega)\leq L_{m-1}(\Lambda(\omega),\omega)=:L_{m-1,\Lambda}(\omega),$
which indicates that $(L_{m,\Lambda},\mathcal{G}_{m})_{m\geq 1}$ is indeed an non-negative super-martingale. Then, combining law of iterated expectation above yields
$\max_{m}\mathbb{E}[L_{m,\Lambda}]\leq 1.$ (A.13)
By letting $m=n$, using with Markov inequality yields
$\mathbb{P}\Big(\Lambda\mathbf{M}_{n}-\frac{\Lambda^{2}}{2}[\mathbf{M}]_{n}\geq t\Big)=\mathbb{P}\Big(\exp\Big(\Lambda\mathbf{M}_{n}-\frac{\Lambda^{2}}{2}[\mathbf{M}]_{n}\Big)\geq e^{t}\Big)=\mathbb{P}(L_{n,\Lambda}\geq e^{t})\leq e^{-t}$ (A.14)
holds for any $t>0$, $n\geq 1$ and any random variable $\Lambda$ taking value in $[0,1/b]$. Step 3 Define
$\mathcal{E}_{t}=\Big\{\Lambda\mathbf{M}_{n}-\frac{\Lambda^{2}}{2}[\mathbf{M}]_{n}\geq t\Big\},\ \ E_{t}=\Big\{\mathbf{M}_{n}\geq\sqrt{2t([\mathbf{M}]_{n}\lor 2b^{2}t)}\Big\},\ \Lambda=\min\Big\{\sqrt{\frac{2t}{[\mathbf{M}]_{n}}},\frac{1}{b}\Big\}.$
Obviously, we can decompose event $E_{t}$ as
$E_{t}=E_{t}^{1}\cup E^{2}_{t},$
where $E_{t}^{1}=E_{t}\cap\{[\mathbf{M}]_{n}\geq 2b^{2}t\}$ and $E_{t}^{2}=E_{t}\cap\{[\mathbf{M}]_{n}<2b^{2}t\}$.
When $\omega\in E^{1}_{t}$, $\Lambda(\omega)=\sqrt{\frac{2t}{[\mathbf{M}]_{n}(\omega)}}$, $\mathbf{M}_{n}(\omega)\geq\sqrt{2[\mathbf{M}]_{n}(\omega)t}$ and $\mathbf{M}]_{n}(\omega)\geq 2b^{2}t$ hold simultaneously. Some basic algebra shows
$\Lambda(\omega)\mathbf{M}_{n}(\omega)-\frac{\Lambda^{2}(\omega)}{2}[\mathbf{M}]_{n}(\omega)\geq\Lambda(\omega)\sqrt{2[\mathbf{M}]_{n}(\omega)t}-\frac{\Lambda^{2}(\omega)}{2}[\mathbf{M}]_{n}(\omega)=2t-t=t,$
where the second last equality is based on the fact that $\Lambda(\omega)=\sqrt{\frac{2t}{[\mathbf{M}]_{n}(\omega)}}$. Thus, for any $t>0$,
$E^{1}_{t}\subset\mathcal{E}_{t}.$ (A.15)
When $\omega\in E^{2}_{t}$, since $\Lambda(\omega)=\frac{1}{b}$, $\mathbf{M}_{n}(\omega)\geq 2bt$ and $\mathbf{M}]_{n}(\omega)<2b^{2}t$ hold simultaneously and
$\Lambda(\omega)\mathbf{M}_{n}(\omega)-\frac{\Lambda^{2}(\omega)}{2}[\mathbf{M}]_{n}(\omega)\geq\frac{1}{b}2bt-\frac{1}{2b^{2}}2b^{2}t=2t-t=t,$
we obtain
$E^{2}_{t}\subset\mathcal{E}_{t}.$ (A.16)
Above all, combining (A.15) and (A.16) yields
$\mathbb{P}(E_{t})\leq\mathbb{P}(\mathcal{E}_{t})\leq e^{-t},\ \forall\ t>0.$
Q.E.D
Proof of Corollary 1 Since $(Z_{i})_{i\geq 0}$ is adapted to $(\mathcal{A}_{i})_{i\geq 0}$ and $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$, $(Z_{i}-\mu,\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence. Then, applying Theorem 1 indicates that, for any given $\delta>0$ and $n\geq 1$, the following event happens with probability no less than $1-2\delta$,
@@
$\left\{\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}\lor\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}=:E.$ (A.17)
@@
Please note that, for any given $a,b,c,d\geq 0$ and $d\leq a$, event $\{a\leq\max\{b+d,c\}\}\subset\{a-d\leq b\lor c\}$. Thus,
$E_{\delta}\subset\left\{\Big(1-\sqrt{2\frac{1}{n}\log(\frac{1}{\delta})}\Big)\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}}{n^{2}}}\lor\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}=E^{\prime}.$ (A.19)
Thus, we obtain $\mathbb{P}(E^{\prime})\geq P(E)\geq 1-2\delta$, which finish the proof.
Proof of Corollary 2 Since the proof of Corollary 2 is a direct applications of Theorem 1 and Corollary 2, we only show the key points here. Since $(Z_{i}-\mu,\mathcal{A}_{i})_{i\geq 0}$ is still a martingale difference sequence, According to (A.17), we construct event
$\mathcal{E}=\left\{\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}<\frac{2b_{V}\log(\frac{1}{\delta})}{n}\right\}.$ (A.20)
@@
$\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}=\sqrt{\frac{2\log(\frac{1}{\delta})[\frac{1}{n}\sum_{i=1}^{n}((Z_{i}-\mu)^{2}-\text{var}(Z_{i}))]}{n}+\frac{2\log(\frac{1}{\delta})\text{Var}(Z_{0})}{n}},$
@@
$\mathcal{E}=\left\{\frac{1}{n}\sum_{i=1}^{n}((Z_{i}-\mu)^{2}-\text{var}(Z_{i}))+\text{Var}(Z_{0})<\frac{2\log(\frac{1}{\delta})b_{V}^{2}}{n}\right\}.$
For any given $\delta>0$ and $\eta\in(0,1)$, $N(\delta,\eta)=\min\{n\geq 1:\eta\text{Var}(Z_{0})\geq\frac{2\log(\frac{1}{\delta})b_{V}^{2}}{n}\}$. Then, for any $n\geq N(\delta,\eta)$, we have
@@
$\Rightarrow$ $\mathbb{P}(E\cap\mathcal{E}^{c})\geq\mathbb{P}(E)-\mathbb{P}(\mathcal{E}^{\prime})\geq 1-2\delta-\mathbb{P}(\mathcal{E}^{\prime})$ (A.21)
Since $\mathbb{E}[(Z_{i}-\mu)^{2}|\mathcal{A}_{i-1}]=\sigma^{2}>0$, using Freedman inequality yields
@@
Finally, combining (A.2), (A.22) and the calculation of (A.18) finishes the proof.
@@
Step 4 (Application of Theorem 1) According to Theorem 1, we can show that, for any $\delta>0$ and arbitrary $\xi_{n}\searrow 0$, the following inequality holds with probability no less than $1-2\delta$,
$|S_{nb_{m}}(h_{n})|\leq\max\left\{\sqrt{2\sum_{j=1}^{m}M_{nj}^{2}\log(\frac{1}{\delta})},2\max_{1\leq j\leq m}|M_{nj}|\log(\frac{1}{\delta})\right\}\leq\max\{A,B\},$
@@
$\left(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{[n/[l]]}}\right)\Big|\frac{1}{n}S_{nb_{m}}(h_{n})\Big|\leq\max\{A^{\prime},B^{\prime}\},$
@@
$B^{\prime}$ $=\frac{2([l](b_{n}-a_{n})+\triangle_{n}\Psi_{n})}{n}.$
Meanwhile, regarding $|\frac{1}{n}r_{n}|\leq\frac{[l]}{n}(b_{n}-a_{n})$,
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)\max\{A^{\prime},B^{\prime}\}+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-2\delta,$
where $\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{[n/[l]]}}\Big)^{-1}$.
@@
Step 1 (Crucial concentration phenomenon) Based on the $r_{n}=\sum_{i=m[l]+1}^{n}(h_{n}(Z_{ni})-\mu)$ introduced in (A.28) and the facts that $\mathbb{E}[r_{n}]=0$ and $\frac{r_{n}}{n-m[l]}\in[a_{n},b_{n}]$, for any $\tau_{1n}\nearrow\infty$ and $t>0$, we have
$\mathbb{P}\Big(\Big|\frac{1}{\tau_{1n}}r_{n}\Big|>t\Big)=\mathbb{P}\Big(\Big|\frac{1}{n-m[l]}r_{n}\Big|>\frac{\tau_{1n}}{(n-m[l])}t\Big)\leq 2\exp\Big(-2\Big(\frac{(\tau_{1n}/[l])t}{(b_{n}-a_{n})}\Big)^{2}\Big),$ (A.36)
where $C_{n}=\frac{n}{n-m[l]}\geq\frac{n}{[l]}$. Similarly, according to the definition of $T_{nj}$ and $R_{nj}$ defined in Step 3 of the proof of Theorems 2 and 3, we have $\mathbb{E}[T_{nj}]=\mathbb{E}[R_{nj}]=0$ and $|T_{nj}|\leq[l](b_{n}-a_{n})$, $|R_{nj}|\leq 2\triangle_{n}\Psi_{n}$. Thus, for any given $\tau_{2n}\nearrow\infty$ and $t>0$, we have

### v3

@@ A.2 Proof of Results in Section 2
Proof of Theorem 1
Step 0 (Brief of strategy) Obviously, for both (8) and (9), we only need to prove one-sided results. Our proof strategy is simple. Based on a method relying on moment generating function with randomized parameter, we first prove the weaker (8). Meanwhile, note that, by denoting $V_{i}:=\mathbb{E}[Z^{2}_{i}|\mathcal{A}_{i-1}]-Z_{i}^{2}$, $V_{i}$ is measurable with respect to $\mathcal{A}_{i}$ and $\mathbb{E}[V_{i}|\mathcal{A}_{i-1}]=0$, which makes $(V_{i},\mathcal{A}_{i})_{i\geq 0}$ a martingale difference sequence as well. Then, by applying (8) again, we can give a high probability bound of the difference between $\langle\mathbf{M}\rangle_{n}$ and $\mathbf{M}]_{n}$ as well. Finally, combining these results together finishes the proof of (9).
Step 1 (Fundamental tools) In this step, we discuss three simple but important mathematical statements which serve as cornerstones of the later steps.
  • S1
    For every given $i\leq n$ and $\lambda\in[0,3/b)$, we have
    $\mathbb{E}[e^{\lambda Z_{i}}|\mathcal{A}_{i-1}]\leq\exp{(\psi(\lambda)\mathbb{E}[Z_{i}^{2}|\mathcal{A}_{i-1}])},\ \ \psi(\lambda)=\frac{\lambda^{2}}{2(1-\frac{\lambda b}{3})},$ (A.9)
    which follows from the standard Bernstein condition for bounded zero-mean variables (e.g. ? (?)).
  • S2
    For any $\nu>0$ and $x\geq 0$, based on $\psi$ introduced in (A.9),
    $\inf_{0\leq\lambda<3/b}\Big\{\frac{\psi(\lambda)}{\lambda}\nu+\frac{x}{\lambda}\Big\}=\sqrt{2\nu x}+\frac{b}{3}x,$ (A.10)
    where the minimum point is obtained at $\lambda(\nu,x):=\frac{\sqrt{2x/\nu}}{1+\frac{b}{3}\sqrt{2x/\nu}}\in(0,3/b)$.
  • S3
    Suppose $\Lambda$ is a random variable taking value in $[0,3/b)$. Based on the function $\psi$ defined in (A.9), for any given $1\leq i\leq n$, we have
    $\mathbb{E}[\exp(\Lambda Z_{i}-\psi(\Lambda)E[Z_{i}^{2}|\mathcal{A}_{i-1}])|\sigma(\Lambda)\lor\mathcal{A}_{i-1}]\leq 1\ \ \text{a.s.}.$ (A.11)
Since the proof of S2 is direct, we omit it here and focus solely on the proof of S3. By denoting $h_{i}(\lambda,\cdot)=\exp(\lambda Z_{i}(\cdot)-\psi(\lambda)\mathbb{E}[Z_{i}^{2}|\mathcal{A}](\cdot))$, S1 indicates that, for each $i\leq n$,
$g_{i}(\lambda,\omega):=\mathbb{E}[h_{i}(\lambda,\cdot)|\mathcal{A}_{i-1}](\omega)\leq 1\ \ \text{a.s. for}\ \ \forall\ \lambda\in[0,3/b).$ (A.12)
Then, since $\Lambda$ has support $[0,3/b)$, using Lemma A.4 and (A.12) yields that, for any $i\leq n$,
[…]
Thus, we finish the proof of S3.
Step 2 (Proof of (8)) We first focus on event
$\mathbf{A}:=\Big\{\mathbf{M}_{n}\geq\sqrt{2\langle\mathbf{M}\rangle_{n}t}+\frac{b}{3}t\Big\}.$
S3 in Step 1 implies that, for any given $\nu\geq 0$ and $t>0$, there exists $\lambda(\nu,t)\in(0,3/b)$ such that
$\frac{\psi(\lambda(\nu,t))}{\lambda(\nu,t)}v+\frac{t}{\lambda(\nu,t)}=\sqrt{2\nu t}+\frac{b}{3}t.$
By letting $\nu=\langle\mathbf{M}\rangle_{n}$ and $\Lambda_{n}=\lambda(\langle\mathbf{M}\rangle_{n},t)$, we obtain
$\mathbf{A}=\Big\{\Lambda_{n}\mathbf{M}_{n}-\psi(\Lambda_{n})\langle\mathbf{M}\rangle_{n}\geq t\Big\}=\Big\{\exp\Big(\Lambda_{n}\mathbf{M}_{n}-\psi(\Lambda_{n})\langle\mathbf{M}\rangle_{n}\Big)\geq e^{t}\Big\}=:\mathbf{B}.$
Then, using Markov inequality yields
$\mathbb{P}\Big(\mathbf{M}_{n}\geq\sqrt{2\langle\mathbf{M}\rangle_{n}t}+\frac{b}{3}t\Big)\leq e^{-t}\mathbb{E}\Big[\exp\Big(\Lambda_{n}\mathbf{M}_{n}-\psi(\Lambda_{n})\langle\mathbf{M}\rangle_{n}\Big)\Big].$
Define
$\mathcal{G}_{t}=\sigma(\Lambda_{n})\lor\mathcal{A}_{t}\ \text{and}\ L_{n,t}=\exp\Big(\Lambda_{n}\mathbf{M}_{t}-\psi(\Lambda_{n})\langle\mathbf{M}\rangle_{t}\Big).$
Obviously, $\{L_{n,t}\}_{t=1}^{n}$ is adapted to filter $\{\mathcal{G}_{t}\}_{t=1}^{n}$. i.e. $L_{n,t}$ is measurable with respect to $\mathcal{G}_{t}$. Meanwhile, S3 of Step 1 asserts that, for any $2\leq t\leq n$
$\mathbb{E}[L_{n,t}|\mathcal{G}_{t-1}]=L_{n,t-1}\mathbb{E}[\exp(\Lambda Z_{i}-\psi(\Lambda)E[Z_{i}^{2}|\mathcal{A}_{i-1}])|\mathcal{G}_{t-1}]\leq L_{n,t-1}\ \ \text{a.s.}.$ (A.14)
Thus, $(L_{n,t},\mathcal{G}_{t})$ is a non-negative super-martingale and $\mathbb{E}[L_{n,n}]\leq\mathbb{E}[L_{n,n-1}]\leq\dots\leq\mathbb{E}[L_{n,1}]\leq 1$. Then, we finish the proof of (8). i.e.
$\mathbb{P}(\mathbf{E}_{1})\geq 1-2e^{-t},\ \ \mathbf{E}_{1}:=\Big\{|\mathbf{M}_{n}|\leq\sqrt{2\langle\mathbf{M}\rangle_{n}t}+\frac{b}{3}t\Big\}.$ (A.15)
Step 3 (Comparing $\mathbf{M}]_{n}$ with $\langle\mathbf{M}\rangle_{n}$ by (8)) As pointed out in Step 0, we can regard $(V_{i},\mathcal{A}_{i})_{i\geq 1}$ as another martingale difference sequence satisfying
  • (A1)
    $V_{i}=\mathbb{E}[Z_{i}^{2}|\mathcal{A}_{i-1}]-Z_{i}^{2}$ and $V_{i}||_{\infty}\leq b^{2}$.
  • (A2)
    $\sum_{i=1}^{n}\mathbb{E}[V_{i}^{2}|\mathcal{A}_{i-1}]\leq b^{2}\sum_{i=1}^{n}E[Z^{2}_{i}|\mathcal{A}_{i-1}]=b^{2}\langle\mathbf{M}\rangle_{n}$.
The reason for point (A2) is because, for each $i$, we have
$\mathbb{E}[V_{i}^{2}|\mathcal{A}_{i-1}]=\mathbb{E}[(Z_{i}^{2}-\mathbb{E}[Z_{i}^{2}|\mathcal{A}_{i-1}])^{2}|\mathcal{A}_{i-1}]=\text{Var}(Z_{i}^{2}|\mathcal{A}_{i-1})\leq\mathbb{E}[Z^{4}_{i}|\mathcal{A}_{i-1}]\leq b^{2}\mathbb{E}[Z_{i}^{2}|\mathcal{A}_{i-1}].$
Since $\langle\mathbf{M}\rangle_{n}-[\mathbf{M}]_{n}=\sum_{i=1}^{n}V_{i}$, together with (A1) and (A2) above, using one-sided version of (8) yields
@@
$\mathbb{P}\left(\langle\mathbf{M}\rangle_{n}-[\mathbf{M}]_{n}\geq\sqrt{2b^{2}\langle\mathbf{M}\rangle_{n}t}+\frac{b^{2}}{3}t\right)\leq e^{-t},$ (A.16)
which indicates
$\mathbb{P}(\mathbf{E}_{2})\geq 1-e^{-t},\ \mathbf{E}_{2}:=\Big\{\langle\mathbf{M}\rangle_{n}\leq[\mathbf{M}]_{n}+\sqrt{2b^{2}\langle\mathbf{M}\rangle_{n}t}+\frac{b^{2}}{3}t\Big\}.$ (A.17)
For each $n$, by denoting $A=[\mathbf{M}]_{n}$, $B=\sqrt{2b^{2}t}$ and $C=\frac{b^{2}}{3}t$ and regarding $\sqrt{\langle\mathbf{M}\rangle_{n}}$ as a non-negative $\mathbb{R}$-valued random variable $\mathbf{Y}$, we have
$\mathbf{E}_{2}=\{\mathbf{Y}^{2}\leq A+B\mathbf{Y}+C\}=\{\mathbf{Y}^{2}-B\mathbf{Y}-(A+C)\leq 0\}=:\{f(\mathbf{Y})\leq 0\}.$
Note that $\mathbf{Y}\geq 0$, $B\geq 0$ and $A+C>0$ (since $C>0$) hold almost surely and, on $\mathbb{R}^{+}$, equation $f(y)=0$ has unique root $y_{\text{root}}=\frac{B+\sqrt{B^{2}+4(A+C)}}{2}$. Thus,
$\mathbf{E}_{2}\subset\Big\{\mathbf{Y}\leq\frac{B+\sqrt{B^{2}+4(A+C)}}{2}\Big\}.$
Meanwhile, since $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ holds for all $x,y\geq 0$, we obtain
$\frac{B+\sqrt{B^{2}+4(A+C)}}{2}\leq\frac{B+B+2\sqrt{A+C}}{2}=B+\sqrt{A+C},$
which implies
$\mathbf{E}_{2}\subset\{\mathbf{Y}\leq B+\sqrt{A+C}\}=:\mathbf{E}_{3}.$
Substituting the definitions of $A$, $B$ and $C$ yields
$\mathbf{E}_{3}=\{\sqrt{2b^{2}t}\mathbf{Y}\leq\sqrt{2b^{2}t}(B+\sqrt{A+C})\}=\{\sqrt{2b^{2}\langle\mathbf{M}\rangle_{n}t}\leq 2b^{2}t+b\sqrt{2t(A+C)}\}$
$\subset$ $\Big\{\sqrt{2b^{2}\langle\mathbf{M}\rangle_{n}t}\leq 2b^{2}t+b\sqrt{2At}+b\sqrt{2Ct}\Big\}=\Big\{\sqrt{2\langle\mathbf{M}\rangle_{n}t}\leq\sqrt{2[\mathbf{M}]_{n}t}+\Big(2+\sqrt{2/3}\Big)bt\Big\}=:\mathbf{E}_{4}.$
Above all, together with (A.17), we obtain
$\mathbb{P}(\mathbf{E}_{4})\geq\mathbb{P}(\mathbf{E}_{3})\geq\mathbb{P}(\mathbf{E}_{2})\geq 1-e^{-t},\ \ \forall\ t>0.$ (A.18)
Step 4 (Conclusion) Combining (A.15) and (A.18) yields that the probabilistic measure of event $\mathbf{E}_{1}\cap\mathbf{E}_{2}$ is no less than $1-3e^{-t}$. Then, we finish the proof of (9) by noticing
$\mathbf{E}_{1}\cap\mathbf{E}_{4}\subset\Big\{\mathbf{M}_{n}\leq\sqrt{2[\mathbf{M}]_{n}t}+(2+\frac{1}{3}+\sqrt{2/3})bt\Big\}$
and $2+\frac{1}{3}+\sqrt{2/3}\leq 3.15$. Thus, we finish the proof of Theorem 1.
Proof of Corollary 1 Since $(Z_{i})_{i\geq 0}$ is adapted to $(\mathcal{A}_{i})_{i\geq 0}$ and $\mathbb{E}[Z_{i}|\mathcal{A}_{i-1}]=\mu$, $(Z_{i}-\mu,\mathcal{A}_{i})_{i\geq 0}$ is a martingale difference sequence. Then, applying Theorem 1 indicates that, for any given $\delta>0$ and $n\geq 1$, the following event happens with probability no less than $1-3\delta$,
$\left\{\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}+\frac{3.15b\log(\frac{1}{\delta})}{n}\right\}=:E.$ (A.19)
@@
Thus, some basic algebra shows
$E_{\delta}\subset\left\{\Big(1-\sqrt{2\frac{1}{n}\log(\frac{1}{\delta})}\Big)\Big|\overline{\mathbf{M}}_{n}-\mu\Big|\leq\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\overline{\mathbf{M}}_{n})^{2}}{n^{2}}}+\frac{3.15b\log(\frac{1}{\delta})}{n}\right\}=E^{\prime}.$ (A.21)
Thus, we obtain $\mathbb{P}(E^{\prime})\geq P(E)\geq 1-3\delta$. Then, we finish the proof by doing the following change of variable, $3\delta=2\alpha$
Proof of Corollary 2 Since the proof of Corollary 2 is a direct applications of Theorem 1 and Corollary 1, we only show the key points here. Since $(Z_{i}-\mu,\mathcal{A}_{i})_{i\geq 0}$ is still a martingale difference sequence, According to (A.19), for some user-defined $\xi_{n}=o(\frac{1}{\sqrt{n}})$, we construct event
$\mathcal{E}=\left\{\xi_{n}\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}<\frac{2b\log(\frac{1}{\delta})}{n}\right\}.$ (A.22)
@@
$\xi_{n}\sqrt{\frac{2\log(\frac{1}{\delta})\sum_{i=1}^{n}(Z_{i}-\mu)^{2}}{n^{2}}}=\xi_{n}\sqrt{\frac{2\log(\frac{1}{\delta})[\frac{1}{n}\sum_{i=1}^{n}((Z_{i}-\mu)^{2}-\text{var}(Z_{i}))]}{n}+\frac{2\log(\frac{1}{\delta})\text{Var}(Z_{0})}{n}},$
@@
$\mathcal{E}=\left\{\frac{1}{n}\sum_{i=1}^{n}((Z_{i}-\mu)^{2}-\text{var}(Z_{i}))+\text{Var}(Z_{0})<\frac{5b^{2}\log(\frac{1}{\delta})}{n\xi_{n}^{2}}\right\}.$
For any given $\delta>0$ and user-defined $\eta\in(0,1)$, define
$N(\delta,\eta)=\min\{n\geq 1:\eta\text{Var}(Z_{0})\geq\frac{5b^{2}\log(\frac{1}{\delta})}{n\xi_{n}^{2}}\}.$
Then, for any $n\geq N(\delta,\eta)$, we have
@@
$\Rightarrow$ $\mathbb{P}(E\cap\mathcal{E}^{c})\geq\mathbb{P}(E)-\mathbb{P}(\mathcal{E}^{\prime})\geq 1-3\delta-\mathbb{P}(\mathcal{E}^{\prime})$ (A.23)
Since $\mathbb{E}[(Z_{i}-\mu)^{2}|\mathcal{A}_{i-1}]=\sigma^{2}>0$ and $Z_{i}$’s are actually IID data, using Bernstein inequality yields
@@
Finally, by combining (A.2), (A.24) and the change of variable, $3\delta=2\alpha$, we finish the proof.
@@
Step 4 (Application of Theorem 1) According to Theorem 1, we can show that, for any $\delta>0$ and arbitrary $\xi_{n}\searrow 0$, the following inequality holds with probability no less than $1-3\delta$,
[…]
@@
$\left(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{[n/[l]]}}\right)\Big|\frac{1}{n}S_{nb_{m}}(h_{n})\Big|\leq A^{\prime}+B^{\prime}$
@@
$B^{\prime}$ $=\frac{3.15([l](b_{n}-a_{n})+\triangle_{n}\Psi_{n})\log(\frac{1}{\delta})}{n}.$
Meanwhile, regarding $|\frac{1}{n}r_{n}|\leq\frac{[l]}{n}(b_{n}-a_{n})$, we obtain
$\mathbb{P}\left(\Big|\frac{1}{n}S_{nn}(h_{n})\Big|\leq\tilde{\nu}_{n}(\delta)(A^{\prime}+B^{\prime})+\frac{[l]}{n}(b_{n}-a_{n})\right)\geq 1-3\delta,$
where $\tilde{\nu}_{n}(\delta)=\Big(1-\sqrt{\frac{2\log(\frac{1}{\delta})}{[n/[l]]}}\Big)^{-1}$. We also finish the proof by letting $3\delta=2\alpha$.
@@
Step 1 (Crucial concentration phenomenon) Based on the $r_{n}=\sum_{i=m[l]+1}^{n}(h_{n}(Z_{ni})-\mu)$ introduced in (A.30) and $\mathbb{E}[r_{n}]=0$. $\frac{r_{n}}{n-m[l]}\in[a_{n},b_{n}]$, for any $\tau_{1n}\nearrow\infty$ and $t>0$, $n\geq 1$, Lemma A.7 yields
[…]
Similarly, according to the definition of $T_{nj}$ and $R_{nj}$ defined in Step 3 of the proof of Theorems 2 and 3, we have $\mathbb{E}[R_{nj}]=0$ and $R_{nj}||_{\infty}\leq 2\triangle_{n}\Psi_{n}$. Thus, for any given $\tau_{2n}\nearrow\infty$ and $t>0$, we have